A hospital director is told that 35% of the treated patients are uninsured. The director wants to test the claim that the percentage of uninsured patients is not the expected percentage. A sample of 120 patients found that 30 were uninsured. At the 0.05 level, is there enough evidence to support the director's claim?

Step 4 of 7 : Determine the P-value of the test statistic. Round your answer to four decimal places.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p = 0.35

Ha : p 0.35

= x / n = 30 /120 = 0.25

1 - P0 = 1 - 0.35 = 0.65

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.25 - 0.35 / [(0.35 * 0.65) / 120]

Test statistic = -2.30

P(z < -2.30) = 2 * 0.0107 = 0.0214

P-value = 0.0214

= 0.05

P-value <

Reject the null hypothesis .

There is sufficient evidence to support directors claim.